a 20% solution of fertilizer is to be mixed with a 80% solution of fertilizer in order to get 120 gallons of a 60% solution. how many gallons of the 20% solution and 80% solution should be mixed?
Answer by checkley77(12844)

.20x+.80*120=.60(120+x)
.20x+96=72+.60x
.20x-.60x=72-96
-.40x=-24
x=-24/-.40
x=60 gallons of 20% solution is used.
Proof:
.20*60+.80*120=.60(120+60)
12+96=.60*180
108=108